OpenStudy (anonymous):

Help maybe? I think I know how to do it, just need reassurance... R, S, and T are collinear. S is between R and T. RS = 2w + 1, ST = w -1, and RT = 18. Use the Segment Addition Postulate to solve for w. Then determine the length of RS.

OpenStudy (anonymous):



OpenStudy (anonymous):

I think I have to add the two equations together? 2w-1 + w-1 = 18

OpenStudy (anonymous):

Does someone care to answer whether or not this is right?

sammixboo (sammixboo):

Yes you have to add the two equations together, because RS + ST = RT

sammixboo (sammixboo):

I believe so

OpenStudy (anonymous):

Ok, and when I do so I get 20/3, which comes out to some sort of decimal...

OpenStudy (anonymous):

However, I dont think this is correct.... :/

OpenStudy (jdoe0001):

RT = RS + ST RT = 18 thus RT = 2w-1 + w-1 18 = 2w-1 + w-1 <--- yeap, you're correct

sammixboo (sammixboo):

20/3 = w? You can check by replacing w with 20/3 and solving it =)

OpenStudy (jdoe0001):

well... you typed in 2w-1, it should be 2w+1 though

OpenStudy (anonymous):

And then do I plug in the answer to the RS equation?

OpenStudy (anonymous):

Yeah... I see the mistake now...

OpenStudy (jdoe0001):

RT = RS + ST RT = 18 thus RT = 2w+1 + w-1 18 = 2w+1 + w-1 <--- there

sammixboo (sammixboo):

Yes, you can if you want

sammixboo (sammixboo):

To see if both sides of the equation is equal

OpenStudy (anonymous):

Thank you so much! TO the both of you @sammixboo and @jdoe0001

sammixboo (sammixboo):

So is this true? 18 = 2*(20/3) + 1 + (20/3) - 1

OpenStudy (jdoe0001):

yw

sammixboo (sammixboo):

No problem =)